### v2

# LBL: Logarithmic Barrier Loss Function for One-class Classification

Abstract
One-class classification (OCC) aims to train a classifier solely on target data and attracts increasing attention due to its applicability in practice. Despite OCC has obtained many advances, it still lacks the effective OCC loss functions for deep learning. In this paper, a novel logarithmic barrier function based OCC loss (LBL) that assigns large gradients to margin samples and thus derives more compact hypersphere is first proposed by approximating the OCC objective smoothly. But the optimization of LBL may be instability especially when samples lie on the boundary leading to the infinity value. To address this issue, a smoother LBLSig loss is further proposed by utilizing a unilateral relaxation Sigmoid function. Experiments on different networks demonstrate the effectiveness of the proposed LBL and LBLSig. The source code can be found at https://github.com/ML-HDU/LBL_LBLSig.
Index Terms—  One-class classification, anomaly detection, logarithmic barrier loss, hypersphere optimization
1 Introduction
One-class classification (OCC), also known as the anomaly detection (AD), aims to construct the model only exploiting the characteristics of target data. It has attracted much attention in various communities such as data mining, machine learning and computer vision [1]. The classical methods [2, 3] are generally required feature engineering. In comparisons, deep learning-based OCC algorithms achieve promising performance including the autoencoder (AE) based methods [4, 5, 6, 7, 8], the generative adversarial network (GAN) based methods [9, 10, 11], and other discrimination-based methods [12, 13, 14, 15, 16].
However, little attention is paid to the OCC loss function design. Recently typical OCC loss functions are developed in [13] and [14]. Ruff et al. [13] proposes the soft-boundary loss (SBL) that can be regarded as an extension of the mean square error (MSE) in OCC. But SBL is only demonstrated on the LeNet-type convolution neural network (CNN) and the performance is not promising. In addition, SBL is essentially the hard sample mining with hinge function form that is non-smooth and difficult to optimize. Different from the MSE-type SBL, a cross entropy (CE)-type loss function named HRN that consists of a negative log-likelihood (NLL) term and a gradient penalty term with respect to inputs is proposed in [14]. The HRN shows the surprising results but the performance both in theory and experiment is only verified on the simple multilayer perceptron (MLP). Particularly, we find that the HRN is optimized difficultly in practice and the performance is poor on CNN. Therefore, it still lacks effective OCC loss functions for deep learning.
Actually, OCC aims to seek a hypersphere with the minimum radius to enclose the target data. In this paper, a novel logarithmic barrier function [17] based OCC loss function (LBL) is first proposed by approximating the OCC objective smoothly. LBL assigns larger gradients to these samples close to boundary, and thus achieves more compact hypersphere. However, the direct approximation by the logarithmic barrier function may cause the instability especially when samples lie on the boundary leading to the infinity loss. To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is further proposed. Owing to the unilateral relaxation Sigmoid function, the optimization becomes smoother. Particularly, LBLSig can be seen as the fusion of MSE and CE. Experiments on various networks are conducted to verify the effectiveness of the proposed LBL/LBLSig. The results demonstrate the superior performance of our proposed algorithms.
2 Problem Statement
The most common MSE-based OCC loss (MSE-OCL) is a simple but effective OCC loss function [18, 19, 8]. Given an OCC training set with $N$ samples ${\bf{x}}_{i}$}, MSE-OCL seeks a hypersphere centered at $\bf{c}$, that is
$\min_{\boldsymbol{\mathcal{W}}}\quad\frac{1}{N}\sum_{i=1}^{N}\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}+\frac{{\lambda}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2},$ (1)
where $\bf{c}$} is the center of hypersphere, $\lambda$ is the trade-off parameter, and ${\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)$ is the model output with the learnable parameter ${\boldsymbol{\mathcal{W}}}=\{{{\bf{W}}^{(1)},\cdots,{\bf{W}}^{(L)}}\}$. It can be readily seen that (1) is a direct extension of MSE-based regression. In addition, the minimum-radius is obtained implicitly in MSE-OCL by minimizing the distance between the outputs of model ${\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)$ and the center $\bf{c}$. Different from that, SBL [13] explicitly minimizes the radius of the hypersphere
$\min_{R,{\boldsymbol{\mathcal{W}}}}$ $R^{2}+{\frac{{\lambda}_{1}}{N}}\sum_{i=1}^{N}\max\left\{{0,\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}-R^{2}}\right\}$ (2)
$+\frac{{\lambda}_{2}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2},$
where $\lambda_{1}$ and $\lambda_{2}$ are the trade-off parameters, $R$ is the radius of hypersphere. It is worth pointing out that the radius is updated in each mini-batch [6] by choosing a certain quantile from the sequence $\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}$, as shown in the implementation of SBL [^1] Code of Deep SVDD: https://github.com/lukasruff/Deep-SVDD-PyTorch.. Therefore, SBL is the iterative hard sample mining in essence.
In contrast to the aforementioned MSE-type OCC losses, HRN [14] can be expressed as
$\small\min_{\boldsymbol{\mathcal{W}}}\quad\underbrace{-\sum_{i=1}^{N}{{\log\left({{\text{Sig}}\left({{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}\right)}\right)}}}_{\text{NLL}}+{\lambda}\underbrace{\sum_{i=1}^{N}{\left\|{{\nabla}_{\bf{x}}{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}}\right\|_{F}^{q}}}_{\text{H-regularization}},$ (3)
where $\lambda$ and $q$ are two manual hyperparameters, and $\text{Sig}(\cdot)$ is the Sigmoid function that can be seen as the probability belonging to targets. It can be readily seen that minimizing NLL means to obtain high output value for target data, and the NLL term is essentially the direct extension from CE. H-regularization term is introduced as a constraint over the growth on the value of ${\boldsymbol{\Phi}}\left({\bf{x}}\right)$} caused by the single target class. HRN shows the superior performance than SBL but it is only verified on MLP. In practice, we find that the HRN is optimized difficultly and the performance is poor on CNN.
3 Proposed methods
3.1 LBL
OCC aims to seek a hypersphere that encloses targets. Therefore, the objective of OCC can be expressed as
$D_{i}=Dis\left({{\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}),{\bf{c}}}\right)\leq R,{\ }i=1,\cdots,N,$ (4)
where $Dis(\cdot)$ represents a distance function such as the L2-norm distance, and $R$ is the radius of the hypersphere. Then, a loss function can be constructed by directly approximating the inequality as
$\min\limits_{\boldsymbol{\mathcal{W}}}{}$ $\frac{1}{N}\sum\limits_{i=1}^{N}{I\left({u_{i}}\right)+\frac{{\lambda}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2}},$ (5)
${\rm{s.}}{\ }{\rm{t.}}{}$ ${u_{i}}={D_{i}^{2}}-R^{2},$
where $I(u)$ is the indicator function that is defined as
$I(u)=\left\{\begin{aligned} 0,{\ }&u\leq 0,\\ \infty,{\ }&u>0.\end{aligned}\right$. (6)
Obviously, minimizing $I(u)$ can enforce ${\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}})$ to fall inside the hypersphere $({\bf{c}},{R})$. However, (5) is difficult to be optimized due to that $I(u)$ is not differentiable. To address this issue, LBL is proposed to approximate (6) smoothly by the logarithmic barrier function [17] as
$\min\limits_{\boldsymbol{\mathcal{W}}}{}$ $-\frac{1}{N}\sum\limits_{i=1}^{N}{\log\left(-{u_{i}}\right)+\frac{{\lambda}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2}},$ (7)
${\rm{s.}}{\ }{\rm{t.}}{}$ ${u_{i}}={D^{2}_{i}}-R^{2}.$
It can be seen that (7) is convex and nondecreasing with respect to $u$. Specially, it increases to $\infty$ smoothly as $u$ increases to 0. In addition, the value of $I(u)$ and its gradient with respect to $u<0$ are both equal to 0 leading to the sample distributions inside the hypersphere unchange. On the contrary, Denote by $J_{1}({\boldsymbol{\mathcal{W}}})=-\frac{1}{N}\sum\limits_{i=1}^{N}{\log\left(-{u_{i}}\right)}$. The gradient with respect to $\boldsymbol{\mathcal{W}}$} can be computed as
${\frac{\partial{J_{1}}}{\partial{\boldsymbol{\mathcal{W}}}}}={\frac{1}{N}}\sum_{i=1}^{N}{{\frac{1}{-u_{i}}}\cdot{\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$ (8)
It can be seen that different penalties are assigned to target samples in LBL. The samples close to the boundary of the hypersphere (${Dis\left({{\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}),{\bf{c}}}\right)}\to R$) will obtain higher penalty as well as the larger gradient value. In this way, LBL may pay more attention to contract these samples to center $\bf{c}$}. Those samples away from the boundary that have small loss and gradient value are fine-tuned in the optimization of LBL for better distributions. However, this advantage may lead to vanishing gradients when fixing hypersphere radius $R$ during training. To address this issue, $R$ is reset to twice the maximum distance when every several epochs. In this way, LBL is optimized by contracting the hypersphere with radius $R$ and thus keeps a fast learning speed. For the hypersphere center $\bf{c}$, it depends on models and will be given in experiments.
3.2 LBLSig
The samples away from the boundary can obtain larger gradients in LBL than those close to center $\bf{c}$, which is similar with SBL but LBL is smoother. However, both SBL and LBL may suffer from the noises that always obtain larger gradients leading to wrong learning. To address this issue, a simple method is to slack the boundary by rejecting a percentage of targets samples, i.e., the radius $R$ is computed by
$R={{Quantile}}\left({Dis\left({{\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}),{\bf{c}}}\right)},q\right),$ (9)
where ${Quantile}(\cdot,q)$ computes the $q$-th quantile of the sequence $\{{Dis\left({{\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}),{\bf{c}}}\right)},i=1,\cdots,N\}$. However, LBL becomes instability especially when samples lie on the boundary and thus fails to solve this situation.
Therefore, a smooth and robust OCC loss function named LBLSig is further developed, which can be expressed as
$\min\limits_{\boldsymbol{\mathcal{W}}}{}$ $-\frac{1}{N}\sum\limits_{i=1}^{N}{\log\left({g(-{u_{i}})}\right)+\frac{{\lambda}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2}},$ (10)
${\rm{s.}}{\ }{\rm{t.}}{}$ ${u_{i}}={D^{2}_{i}}-R^{2}.$
Here,
$g(-u_{i})=\left\{\begin{aligned} {\text{Sig}(-{u_{i}})},{\ }&u_{i}\leq Q,\\ {\text{Sig}(-{Q})},{\ }&u_{i}>Q,\end{aligned}\right$. (11)
where $Q>0$ is a relaxation hyperparameter that controls the tolerance of noisy samples. For $u_{i}>Q$, the Sigmoid function is truncated and the corresponding gradients are equal to 0. Therefore, noisy samples can be filtered. For $u_{i}\leq Q$, denote by $v_{i}={\text{Sig}}(-u_{i})$ and $J_{2}({\boldsymbol{\mathcal{W}}})=-\frac{1}{N}\sum\limits_{i=1}^{N}{\log\left({\text{Sig}(-{u_{i}})}\right)}$. The gradient of $J_{2}({\boldsymbol{\mathcal{W}}})$ with respect to $\boldsymbol{\mathcal{W}}$} can be derived as
${\frac{\partial{J_{2}}}{\partial{\boldsymbol{\mathcal{W}}}}}$ $={\frac{1}{N}}\sum_{i=1}^{N}{\left({1-v_{i}}\right){\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$ (12)
$v_{i}$ can be seen as the probability belonging to targets. Thus, samples close to center have large probability and thus small $\left({1-v_{i}}\right)$, but the samples falling on the boundary and outside have small $v_{i}$ and thus the large gradient value. It is obvious that margin samples need to be assigned larger gradient for optimization than those close to the center $\bf{c}$.
Note that the proposed LBLSig is similar with the NLL term of HRN in (3) but they are different in essence. The probability belonging to targets is obtained in HRN by mapping the model output to $[0,1]$ directly. H-regularization hence has to be used to limit the growth of the model output. Different from that, LBLSig minimizes $Dis\left({{\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}),{\bf{c}}}\right)$} implicitly, which can be observed from (12) that the gradient value relies on $\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}$} but the probability belonging to target $v_{i}$ is introduced to control the gradient. Particularly, LBLSig minimizes MSE-OCL implicitly when using L2-norm. Meanwhile, $J_{2}({\boldsymbol{\mathcal{W}}})$ is a obvious CE-type loss. Thus, LBLSig can be seen as the fusion of MSE and CE.
4 Experiments
4.1 Implementation details
The proposed LBL and LBLSig are conducted on $4$ different networks for comparisons with SBL [13] and HRN [14], including DSLeNet [13], MLP with $2$ hidden layers, and OCITN [15]. For DSLeNet, the center $\bf{c}$ is set to be the mean of the outputs of the initialized DSLeNet on training data. For MLP, the mean of the training samples is computed as the hypersphere center $\bf{c}$. For OCITN, the “Lena image” is chosen as the center $\bf{c}$ the same as [15] (HRN is hence inapplicable to OCITN). The grid search method is employed to select the optimal hyperparameters for all algorithms. More detailed settings can be found in source code due to the page limitation. In addition, SBL argues in [13] that networks with bias will lead to hypersphere collapse. Actually, SBL is non-convex and the network with bias leading to hypersphere collapse is a small probability. We hence do not add any constrains to the bias term in this paper.
Comparisons on $13$ non-image datasets and $1$ image dataset CIFAR10222https://www.cs.toronto.edu/~kriz/cifar.html are conducted. The specification of non-image datasets can be found in [20, 8], and the first class is set as the target and the remaining classes are regarded as outliers. For CIFAR10, each class is respectively chosen as the target class. The complete testing dataset is used to verify the model. The area under the curve (AUC) of the receiver operating characteristic is used as evaluation metric.
Table 1: Comparisons of AUCs with SBL and HRN on CIFAR10.
Backbone Methods Target Class Avg
AIRPLANE AUTOMOBILE BIRD CAT DEER DOG FROG HORSE SHIP TRUCK
DSLeNet SBL [13] 61.7 64.8 49.5 56.0 59.1 62.1 67.8 65.2 75.6 71.0 63.28
HRN 64.73 64.05 55.59 51.81 62.99 50.06 64.13 53.53 71.51 66.8 60.52
LBL 72.51 71.00 54.82 56.79 63.04 60.24 58.86 71.92 78.91 74.27 66.24
LBLSig 72.68 70.26 54.95 58.81 63.52 64.47 70.76 72.25 78.89 74.25 68.08
MLP SBL 76.43 71.60 63.14 63.21 74.55 64.36 76.89 66.93 83.45 75.53 71.61
HRN 58.80 62.27 51.80 54.94 51.01 53.13 56.44 53.02 60.00 63.84 56.53
LBL 76.80 70.72 66.52 63.66 76.63 65.16 80.47 66.57 83.29 77.83 72.77
LBLSig 76.81 71.74 64.06 63.30 75.39 64.98 79.87 66.97 83.41 77.28 72.38
OCITN SBL 66.17 64.32 60.40 54.21 61.00 58.74 66.43 65.06 75.22 68.88 64.04
HRN N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LBL 76.38 87.93 68.82 60.50 71.01 72.17 79.83 80.82 84.33 84.36 76.62
LBLSig 78.03 87.29 69.42 60.91 71.18 71.99 78.54 81.59 84.80 84.96 76.87
Table 3: Comparisons of AUCs with the SOTA OCC algorithms.
Target LSAR OCGAN HLS-OC MOCCA HRN LBL LBLSig
Class [7] [10] [8] [6] [14]
AIRPLANE 73.5 75.7 73.7 66.0 77.3 76.38 78.03
AUTOMOBILE 58.0 53.1 74.4 74.6 69.9 87.93 87.29
BIRD 69.0 64.0 60.9 57.5 60.6 68.82 69.42
CAT 54.2 62.0 63.9 60.1 64.4 60.50 60.91
DEER 76.1 72.3 71.0 61.5 71.5 71.01 71.18
DOG 54.6 62.0 64.1 68.4 67.4 72.17 71.99
FROG 75.1 72.3 78.8 67.4 77.4 79.83 78.54
HORSE 53.5 57.5 70.6 72.1 64.9 80.82 81.59
SHIP 71.7 82.0 81.8 79.2 82.5 84.33 84.80
TRUCK 54.8 55.4 79.0 77.3 77.3 84.36 84.96
AVG 64.05 65.63 71.82 68.61 71.32 76.62 76.87
  • *Conducted on OCITN
4.2 Comparisons with SBL and HRN
Table 1 gives comparisons of AUCs on CIFAR10 where the best results are highlighted in bold font.
DSLeNet as backbone. The results of SBL are taken from [13] for comparisons. The proposed LBLSig obtains the highest AUC value on average. Specially, LBLSig outperforms the SBL [13] on all classes, and generally has superior performance than HRN.
MLP as backbone. LBL and LBLSig obtain the first and second best performance on average, respectively, and much higher performance than HRN in our MLP structure. Particularly, the results in [14] are obtained using much more complicated MLP than ours. Table 3 lists the results in [14] but LBL and LBLSig still obtain better performance than HRN of [14], while only 2-hidden-layer MLP is used by us.
OCITN as backbone. LBLSig performs the best on 7 out of 10 classes and achieves the highest average AUC. It should be pointed out that HRN is hard to be trained on OCITN, and thus the results fails to be given. In other words, LBL and LBLSig have wider applicability than HRN. In addition, comparisons among networks show that LBL/LBLSig obtains the best performance when using OCITN as backbone.
At last, comparisons on non-image datasets are provided in Table 2, where the single hidden layer MLP is used. Obviously, LBL/LBLSig outperform SBL and HRN on all non-image datasets and LBLSig wins the best average AUC value.
4.3 Comparisons with SOTA algorithms
Comparisons on CIFAR10 with common state-of-the-art (SOTA) OCC algorithms, including LSAR [7], OCGAN [10], HLS-OC [8], MOCCA [6] and HRN [14], are reported. OCITN is chosen as the backbone to conduct LBL/LBLSig. Results are listed in Tabel 3 and it can be seen that the best performance is obtained by the proposed LBL/LBLSig. Specially, LBLSig achieves more than $12\%$, $10\%$, $4\%$, $8\%$ and $5\%$ increments of the average AUC over the LSAR, OCGAN, HLS-OC, MOCCA and HRN, respectively.
5 Conclusions
In this paper, two novel OCC loss functions named LBL and LBLSig were proposed. LBL employed the logarithmic barrier function to approximate the OCC objective smoothly. It assigned large gradients to margin samples and thus derived more compact hypersphere. LBLSig was further proposed to improve the optimization smoothness and stability by utilizing a unilateral relaxation Sigmoid function. LBLSig can be seen as the fusion of MSE and CE. The superior performance of the proposed LBL and LBLSig was experimentally verified in comparisons with several state-of-the-art OCC algorithms.
Table 2: Comparisons of AUCs on non-image datasets.
Dataset SBL HRN LBL LBLSig
abalone 87.95 87.69 87.97 88.03
Arrhythmia 76.49 73.50 76.68 77.17
BASEHOCK 57.25 56.47 57.47 59.82
diabetes 72.47 69.77 71.91 73.04
Diabetic 63.36 62.74 65.45 63.68
ecoli 89.09 91.59 96.24 94.58
heart 83.51 82.39 82.72 85.61
leukemia 81.43 69.29 83.69 83.45
liver 58.79 61.25 58.65 61.91
magic 83.51 85.87 86.93 84.36
Online news 61.30 60.35 62.22 62.37
RCV1-4Class 78.27 54.84 79.67 79.05
Sonar 69.79 62.18 70.36 72.73
AVG 74.09 70.61 75.38 75.83

### v3

@@
 Abstract
-One-class classification (OCC) aims to train a classifier solely on target data and attracts increasing attention due to its applicability in practice.
-Despite OCC has obtained many advances, it still lacks the effective OCC loss functions for deep learning.
-In this paper, a novel logarithmic barrier function based OCC loss (LBL) that assigns large gradients to margin samples and thus derives more compact hypersphere is first proposed by approximating the OCC objective smoothly.
-But the optimization of LBL may be instability especially when samples lie on the boundary leading to the infinity value.
-To address this issue, a smoother LBLSig loss is further proposed by utilizing a unilateral relaxation Sigmoid function.
-Experiments on different networks demonstrate the effectiveness of the proposed LBL and LBLSig.
+One-class classification (OCC) aims to train a classifier only with the target class data and attracts great attention for its strong applicability in real-world application.
+Despite a lot of advances have been made in OCC, it still lacks the effective OCC loss functions for deep learning.
+In this paper, a novel logarithmic barrier function based OCC loss (LBL) that assigns large gradients to the margin samples and thus derives more compact hypersphere, is first proposed by approximating the OCC objective smoothly.
+But the optimization of LBL may be instability especially when samples lie on the boundary leading to the infinity loss.
+To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is proposed.
+The LBLSig can be seen as the fusion of the mean square error (MSE) and the cross entropy (CE) and the optimization of LBLSig is smoother owing to the unilateral relaxation Sigmoid function.
+The effectiveness of the proposed LBL and LBLSig is experimentally demonstrated in comparisons with several popular OCC algorithms on different network structures.
 The source code can be found at https://github.com/ML-HDU/LBL_LBLSig.
-Index Terms—  One-class classification, anomaly detection, logarithmic barrier loss, hypersphere optimization
-1 Introduction
-One-class classification (OCC), also known as the anomaly detection (AD), aims to construct the model only exploiting the characteristics of target data.
-It has attracted much attention in various communities such as data mining, machine learning and computer vision [1].
-The classical methods [2, 3] are generally required feature engineering.
-In comparisons, deep learning-based OCC algorithms achieve promising performance including the autoencoder (AE) based methods [4, 5, 6, 7, 8], the generative adversarial network (GAN) based methods [9, 10, 11], and other discrimination-based methods [12, 13, 14, 15, 16].
+I Introduction
+One-class classification (OCC), also known as the anomaly detection (AD), aims to construct the model only exploiting the characteristics of the target data.
+It has attracted much attention in various communities, such as data mining, machine learning and computer vision [17, 36], due to its significance in many applications [14, 30].
+The classical methods, such as the one-class support vector machine (OCSVM) [27] and the support vector data description (SVDD) [28], are confronted with difficulties when dealing with high-dimensional and complex data, and such shallow methods are always required feature engineering.
+In contrast, deep learning-based OCC algorithms achieve exceptional successes, including the autoencoder (AE) based methods [19, 37, 11, 33, 24, 3, 6, 16, 1, 32, 4, 34, 9], the generative adversarial network (GAN) based methods [26, 22, 35, 2, 23, 18, 25], and other discrimination-based methods [10, 29, 21, 13, 12, 20, 7].
 However, little attention is paid to the OCC loss function design.
@@
 Ruff et al.
-[13] proposes the soft-boundary loss (SBL) that can be regarded as an extension of the mean square error (MSE) in OCC.
+[21] propose the soft-boundary loss (SBL) that can be regarded as an extension of the mean square error (MSE) in OCC.
 But SBL is only demonstrated on the LeNet-type convolution neural network (CNN) and the performance is not promising.
-In addition, SBL is essentially the hard sample mining with hinge function form that is non-smooth and difficult to optimize.
+In addition, SBL is essentially the hard sample mining with non-smooth hinge function form.
 Different from the MSE-type SBL, a cross entropy (CE)-type loss function named HRN that consists of a negative log-likelihood (NLL) term and a gradient penalty term with respect to inputs is proposed in [13].
-The HRN shows the surprising results but the performance both in theory and experiment is only verified on the simple multilayer perceptron (MLP).
-Particularly, we find that the HRN is optimized difficultly in practice and the performance is poor on CNN.
-Therefore, it still lacks effective OCC loss functions for deep learning.
-Actually, OCC aims to seek a hypersphere with the minimum radius to enclose the target data.
-In this paper, a novel logarithmic barrier function [17] based OCC loss function (LBL) is first proposed by approximating the OCC objective smoothly.
-LBL assigns larger gradients to these samples close to boundary, and thus achieves more compact hypersphere.
-However, the direct approximation by the logarithmic barrier function may cause the instability especially when samples lie on the boundary leading to the infinity loss.
+OCC is indeed a binary classification obeying the Bernoulli distribution.
+However, the biggest difference from the binary classification is that the non-target samples are absent in training.
+The training of cross entropy solely on target samples may result in malformed learnable parameters.
+The gradient regularization is employed to alleviated this issue, but the training becomes harder.
+In addition, HRN is only verified on the simple multilayer perceptron (MLP).
+In this paper, we follow the OCC design that seeks a hypersphere with the minimum radius to enclose the target data.
+Specifically, a novel logarithmic barrier function based OCC loss function (LBL) is first proposed by approximating the OCC objective smoothly.
+LBL assigns larger gradients to samples close to boundary, and thus achieves more compact hypersphere.
+However, the direct approximation by the logarithmic barrier function may cause the instability when samples lie on the boundary leading to the infinity loss.
 To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is further proposed.
@@
 The results demonstrate the superior performance of our proposed algorithms.
-2 Problem Statement
+II Brief Review
 The most common MSE-based OCC loss (MSE-OCL) is a simple but effective OCC loss function [15, 8, 32].
@@
 where $\lambda_{1}$ and $\lambda_{2}$ are the trade-off parameters, $R$ is the radius of hypersphere.
-It is worth pointing out that the radius is updated in each mini-batch [6] by choosing a certain quantile from the sequence $\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}$, as shown in the implementation of SBL [^1] Code of Deep SVDD: https://github.com/lukasruff/Deep-SVDD-PyTorch..
+It is worth pointing out that the radius is updated in each mini-batch [16] by choosing a certain quantile from the sequence $\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}$, as shown in the implementation [^1] Code of Deep SVDD: https://github.com/lukasruff/Deep-SVDD-PyTorch..
 Therefore, SBL is the iterative hard sample mining in essence.
 In contrast to the aforementioned MSE-type OCC losses, HRN [13] can be expressed as
-$\small\min_{\boldsymbol{\mathcal{W}}}\quad\underbrace{-\sum_{i=1}^{N}{{\log\left({{\text{Sig}}\left({{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}\right)}\right)}}}_{\text{NLL}}+{\lambda}\underbrace{\sum_{i=1}^{N}{\left\|{{\nabla}_{\bf{x}}{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}}\right\|_{F}^{q}}}_{\text{H-regularization}},$ (3)
+$\min_{\boldsymbol{\mathcal{W}}}\quad\underbrace{-\sum_{i=1}^{N}{{\log\left({{\text{Sig}}\left({{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}\right)}\right)}}}_{\text{NLL}}+{\lambda}\underbrace{\sum_{i=1}^{N}{\left\|{{\nabla}_{\bf{x}}{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}}\right\|_{F}^{q}}}_{\text{H-regularization}},$ (3)
 where $\lambda$ and $q$ are two manual hyperparameters, and $\text{Sig}(\cdot)$ is the Sigmoid function that can be seen as the probability belonging to targets.
@@
 H-regularization term is introduced as a constraint over the growth on the value of ${\boldsymbol{\Phi}}\left({\bf{x}}\right)$} caused by the single target class.
-HRN shows the superior performance than SBL but it is only verified on MLP.
-In practice, we find that the HRN is optimized difficultly and the performance is poor on CNN.
-3 Proposed methods
-3.1 LBL
+III Proposed methods
+III-A LBL
 OCC aims to seek a hypersphere that encloses targets. Therefore, the objective of OCC can be expressed as
@@
 However, (5) is difficult to be optimized due to that $I(u)$ is not differentiable.
-To address this issue, LBL is proposed to approximate (6) smoothly by the logarithmic barrier function [17] as
+To address this issue, LBL is proposed to approximate (6) smoothly by the logarithmic barrier function [5, 38] as
 $\min\limits_{\boldsymbol{\mathcal{W}}}{}$ $-\frac{1}{N}\sum\limits_{i=1}^{N}{\log\left(-{u_{i}}\right)+\frac{{\lambda}}{2}\sum_{\ell=1}^{L}\left\|{\bf{W}}^{(\ell)}\right\|_{F}^{2}},$ (7)
@@
 For the hypersphere center $\bf{c}$, it depends on models and will be given in experiments.
-3.2 LBLSig
+III-B LBLSig
 The samples away from the boundary can obtain larger gradients in LBL than those close to center $\bf{c}$, which is similar with SBL but LBL is smoother.
@@
 The gradient of $J_{2}({\boldsymbol{\mathcal{W}}})$ with respect to $\boldsymbol{\mathcal{W}}$} can be derived as
-${\frac{\partial{J_{2}}}{\partial{\boldsymbol{\mathcal{W}}}}}$ $={\frac{1}{N}}\sum_{i=1}^{N}{\left({1-v_{i}}\right){\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$ (12)
+${\frac{\partial{J_{2}}}{\partial{\boldsymbol{\mathcal{W}}}}}$ $=\sum_{i=1}^{N}{{\frac{\partial{J_{2}}}{\partial{v_{i}}}}{\frac{\partial{v_{i}}}{\partial{u_{i}}}}{\frac{\partial{u_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}$ (12)
+$={\frac{1}{N}}\sum_{i=1}^{N}{{\frac{e^{u_{i}}}{1+e^{u_{i}}}}{\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}$
+$={\frac{1}{N}}\sum_{i=1}^{N}{\left({1-v_{i}}\right){\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$
 $v_{i}$ can be seen as the probability belonging to targets.
@@
 Thus, LBLSig can be seen as the fusion of MSE and CE.
-4 Experiments
-4.1 Implementation details
-The proposed LBL and LBLSig are conducted on $4$ different networks for comparisons with SBL [13] and HRN [14], including DSLeNet [13], MLP with $2$ hidden layers, and OCITN [15].
-For DSLeNet, the center $\bf{c}$ is set to be the mean of the outputs of the initialized DSLeNet on training data.
-For MLP, the mean of the training samples is computed as the hypersphere center $\bf{c}$.
-For OCITN, the “Lena image” is chosen as the center $\bf{c}$ the same as [15] (HRN is hence inapplicable to OCITN).
-The grid search method is employed to select the optimal hyperparameters for all algorithms.
-More detailed settings can be found in source code due to the page limitation.
-In addition, SBL argues in [13] that networks with bias will lead to hypersphere collapse.
-Actually, SBL is non-convex and the network with bias leading to hypersphere collapse is a small probability.
-We hence do not add any constrains to the bias term in this paper.
-Comparisons on $13$ non-image datasets and $1$ image dataset CIFAR10222https://www.cs.toronto.edu/~kriz/cifar.html are conducted.
-The specification of non-image datasets can be found in [20, 8], and the first class is set as the target and the remaining classes are regarded as outliers.
-For CIFAR10, each class is respectively chosen as the target class.
-The complete testing dataset is used to verify the model.
-The area under the curve (AUC) of the receiver operating characteristic is used as evaluation metric.
-Table 1: Comparisons of AUCs with SBL and HRN on CIFAR10.
+IV Experiments
+TABLE I: Comparisons of AUCs with SBL and HRN on CIFAR10.
 Backbone Methods Target Class Avg
@@
 LBLSig 78.03 87.29 69.42 60.91 71.18 71.99 78.54 81.59 84.80 84.96 76.87
-Table 3: Comparisons of AUCs with the SOTA OCC algorithms.
-Target LSAR OCGAN HLS-OC MOCCA HRN LBL LBLSig
-Class [7] [10] [8] [6] [14]
-AIRPLANE 73.5 75.7 73.7 66.0 77.3 76.38 78.03
-AUTOMOBILE 58.0 53.1 74.4 74.6 69.9 87.93 87.29
-BIRD 69.0 64.0 60.9 57.5 60.6 68.82 69.42
-CAT 54.2 62.0 63.9 60.1 64.4 60.50 60.91
-DEER 76.1 72.3 71.0 61.5 71.5 71.01 71.18
-DOG 54.6 62.0 64.1 68.4 67.4 72.17 71.99
-FROG 75.1 72.3 78.8 67.4 77.4 79.83 78.54
-HORSE 53.5 57.5 70.6 72.1 64.9 80.82 81.59
-SHIP 71.7 82.0 81.8 79.2 82.5 84.33 84.80
-TRUCK 54.8 55.4 79.0 77.3 77.3 84.36 84.96
-AVG 64.05 65.63 71.82 68.61 71.32 76.62 76.87
-  • *Conducted on OCITN
-4.2 Comparisons with SBL and HRN
-Table 1 gives comparisons of AUCs on CIFAR10 where the best results are highlighted in bold font.
-DSLeNet as backbone.
-The results of SBL are taken from [13] for comparisons.
-The proposed LBLSig obtains the highest AUC value on average.
-Specially, LBLSig outperforms the SBL [13] on all classes, and generally has superior performance than HRN.
-MLP as backbone.
-LBL and LBLSig obtain the first and second best performance on average, respectively, and much higher performance than HRN in our MLP structure.
-Particularly, the results in [14] are obtained using much more complicated MLP than ours.
-Table 3 lists the results in [14] but LBL and LBLSig still obtain better performance than HRN of [14], while only 2-hidden-layer MLP is used by us.
-OCITN as backbone.
-LBLSig performs the best on 7 out of 10 classes and achieves the highest average AUC.
-It should be pointed out that HRN is hard to be trained on OCITN, and thus the results fails to be given.
-In other words, LBL and LBLSig have wider applicability than HRN.
-In addition, comparisons among networks show that LBL/LBLSig obtains the best performance when using OCITN as backbone.
-At last, comparisons on non-image datasets are provided in Table 2, where the single hidden layer MLP is used.
-Obviously, LBL/LBLSig outperform SBL and HRN on all non-image datasets and LBLSig wins the best average AUC value.
-4.3 Comparisons with SOTA algorithms
-Comparisons on CIFAR10 with common state-of-the-art (SOTA) OCC algorithms, including LSAR [7], OCGAN [10], HLS-OC [8], MOCCA [6] and HRN [14], are reported.
-OCITN is chosen as the backbone to conduct LBL/LBLSig.
-Results are listed in Tabel 3 and it can be seen that the best performance is obtained by the proposed LBL/LBLSig.
-Specially, LBLSig achieves more than $12\%$, $10\%$, $4\%$, $8\%$ and $5\%$ increments of the average AUC over the LSAR, OCGAN, HLS-OC, MOCCA and HRN, respectively.
-5 Conclusions
-In this paper, two novel OCC loss functions named LBL and LBLSig were proposed.
-LBL employed the logarithmic barrier function to approximate the OCC objective smoothly.
-It assigned large gradients to margin samples and thus derived more compact hypersphere.
-LBLSig was further proposed to improve the optimization smoothness and stability by utilizing a unilateral relaxation Sigmoid function.
-LBLSig can be seen as the fusion of MSE and CE.
-The superior performance of the proposed LBL and LBLSig was experimentally verified in comparisons with several state-of-the-art OCC algorithms.
-Table 2: Comparisons of AUCs on non-image datasets.
+TABLE II: Comparisons of AUCs on non-image datasets.
 Dataset SBL HRN LBL LBLSig
@@
 AVG 74.09 70.61 75.38 75.83
+IV-A Implementation details
+The proposed LBL and LBLSig are conducted on $4$ different networks for comparisons with SBL [21] and HRN [13], including DSLeNet [21], MLP with $2$ hidden layers, and OCITN [12].
+For DSLeNet, the center $\bf{c}$ is set to be the mean of the outputs of the initialized DSLeNet on training data.
+For MLP, the mean of the training samples is computed as the hypersphere center $\bf{c}$.
+For OCITN, the center $\bf{c}$ is determined the same as [12].
+The grid search method is employed to select the optimal hyperparameters for all algorithms.
+More detailed settings can be found in source code due to the page limitation.
+In addition, SBL argues in [21] that networks with bias will lead to hypersphere collapse.
+Actually, SBL is non-convex and the network with bias leading to hypersphere collapse is a small probability.
+We hence do not add any constrains to the bias term in this paper.
+Comparisons on $13$ non-image datasets and $1$ image dataset CIFAR10222https://www.cs.toronto.edu/~kriz/cifar.html are conducted.
+The specification of non-image datasets can be found in [31, 32], and the first class is set as the target and the remaining classes are regarded as outliers.
+For CIFAR10, each class is respectively chosen as the target class.
+The complete testing dataset is used to verify the model.
+The area under the curve (AUC) of the receiver operating characteristic is used as evaluation metric.
+TABLE III: Comparisons of AUCs with the SOTA OCC algorithms.
+Target LSAR OCGAN HLS-OC MOCCA HRN LBL LBLSig
+Class [1] [18] [32] [16] [13]
+AIRPLANE 73.5 75.7 73.7 66.0 77.3 76.38 78.03
+AUTOMOBILE 58.0 53.1 74.4 74.6 69.9 87.93 87.29
+BIRD 69.0 64.0 60.9 57.5 60.6 68.82 69.42
+CAT 54.2 62.0 63.9 60.1 64.4 60.50 60.91
+DEER 76.1 72.3 71.0 61.5 71.5 71.01 71.18
+DOG 54.6 62.0 64.1 68.4 67.4 72.17 71.99
+FROG 75.1 72.3 78.8 67.4 77.4 79.83 78.54
+HORSE 53.5 57.5 70.6 72.1 64.9 80.82 81.59
+SHIP 71.7 82.0 81.8 79.2 82.5 84.33 84.80
+TRUCK 54.8 55.4 79.0 77.3 77.3 84.36 84.96
+AVG 64.05 65.63 71.82 68.61 71.32 76.62 76.87
+  • *Conducted on OCITN
+IV-B Comparisons with SBL and HRN
+Table I gives the comparisons of AUCs on CIFAR10 under 3 different networks where the best results are highlighted in bold font.
+DSLeNet as backbone.
+The results of SBL are taken from [21] for comparisons.
+The proposed LBLSig obtains the highest AUC value on average.
+Specially, LBLSig outperforms the SBL [21] on all classes and the increments are $10.98\%$ on AIRPLANE, $7.05\%$ on HORSE and $4.8\%$ on average.
+Similarly, the LBLSig generally has superior performance than HRN except the AUTOMOBILE class, and achieves more than $14.41\%$ increments on DOG, $18.72\%$ increments on HORSE and $7.56\%$ increments on average.
+MLP as backbone.
+It can be readily seen that LBL and LBLSig obtain the first and second best performance on average, respectively.
+Particularly, LBL and LBLSig obtain much higher performance than HRN in our MLP structure.
+But it is worthy pointing out that the results reported in [13] are obtained using the structure of $\{3\times[1024$-300]}300]\}-[900[900-300]300]-[300[300-$1]$ which is more complicated than our 2-hidden-layer MLP.
+The results in [13] are listed in Table III, and it can be readily seen by comparisons that LBL and LBLSig generally outperform HRN of [13] and obtain better average AUC while only 2-hidden-layer MLP is used by us.
+OCITN as backbone.
+As shown in Table I, LBLSig performs the best on 7 out of 10 classes and achieves the highest average AUC.
+On the contrary, the SBL on OCITN fails to obtain the promising performance and LBLSig has average $12.83\%$ improvements over SBL.
+In addition, HRN is originally required to fit a scalar, but OCITN aims to a color image, which leads to the hard optimization of HRN on OCITN.
+The results are hence fails to given in the table.
+However, it can be also said that LBL and LBLSig have wider applicability than HRN.
+In addition, the comparisons among networks show that LBL/LBLSig obtains the best performance when using OCITN as backbone.
+Particularly, LBLSig on OCITN achieves $4.49\%$ and $8.79\%$ increments over DSLeNet and MLP, respectively.
+It should be pointed out that LBL may suffer from instability optimization and low robustness and thus LBLSig is developed to improve stability and robustness.
+LBLSig performs a litter better than LBL in experimental comparisons because the experiments are conducted on clean data such as CIFAR10.
+Due to the page limitation, the comparisons of robustness will demonstrated in future work.
+At last, the comparisons on non-image datasets are conducted and shown in Table II, where the single hidden layer MLP is used as the backbone.
+It can be readily seen that LBL/LBLSig outperform SBL and HRN on all non-image datasets and LBLSig wins the best average AUC value.
+IV-C Comparisons with SOTA algorithms
+Comparisons on CIFAR10 with common state-of-the-art (SOTA) OCC algorithms, including LSAR [1], OCGAN [18], HLS-OC [32], MOCCA [16] and HRN [13], are reported.
+OCITN is chosen as the backbone to conduct LBL/LBLSig.
+Results are listed in Tabel III and it can be seen that the best performance is obtained by the proposed LBL/LBLSig.
+Specially, LBLSig achieves more than $12\%$, $10\%$, $4\%$, $8\%$ and $5\%$ increments of the average AUC over the LSAR, OCGAN, HLS-OC, MOCCA and HRN, respectively.
+V Conclusions
+In this paper, two novel OCC loss functions named LBL and LBLSig were proposed.
+LBL employed the logarithmic barrier function to approximate the OCC objective smoothly.
+It assigned large gradients to margin samples and thus derived more compact hypersphere.
+LBLSig was further proposed to improve the optimization smoothness and stability by utilizing a unilateral relaxation Sigmoid function.
+LBLSig can be seen as the fusion of MSE and CE.
+The superior performance of the proposed LBL and LBLSig was experimentally verified in comparisons with several state-of-the-art OCC algorithms.